# Domain Adaptation from Generated Multi-Weather Images for Unsupervised Maritime Object Classification

### v2

@@ 4.4. Effectiveness of Domain Adaptation
Comparison with the basic networks. ViT-small (Dosovitskiy et al., 2021) and ViT-base (Dosovitskiy et al., 2021) have applied Transformer to image classification for the first time, demonstrating the potential of global attention in visual tasks. On this basis, Swin Transformer (Liu et al., 2021) takes into account global modeling and local concentration bias through windowing and hierarchical design. Compared with these basic networks, our proposed method gained of over 10 % on the classification accuracy, which attributes to our effective designs based on a series of modules in order to mitigate domain divergence.
@@
Comparison with methods using pseudo labels. CDTrans (Xu et al., 2021) learned domain-invariant features through cross-domain attention and used target domain pseudo labels for self-supervised learning. Given the substantial differences in distribution between AIMO and RMO datasets, directly aligning image pairs from the source and target domains is not an effective approach. And the large domain gap makes the pseudo labels more unreliable, which directly affects the model performance. Therefore, CDTrans (Xu et al., 2021) exhibits suboptimal performance. Also, different from PMTrans (Zhu et al., 2023), which constructed intermediate domains to alleviate domain gap and also relied on the pseudo labels during training, our method utilizes VLMs to alleviate the domain-shift from source domain to target domain, achieving nearly 7.85 % gains on the accuracy.
@@
Comparison with method using adversarial adaptation. TVT (Yang et al., 2023) designed a transferability adaption module based on adversarial adaptation. The learned transferabilities were injected into attention blocks, promoting ViT to focus on both domain-invariant and domain-specific features. Compared with TVT (Yang et al., 2023), which also utilizes adversarial training for UDA, the accuracy of our proposed method is 10 % higher, probably due to the use of VLMs (i.e., CLIP) for improving robustness of the source domain features.
Comparison with method using VLMs. PDA (Bai et al., 2024) performed a two-branch prompt tuning. Firstly, the zero-shot capability of CLIP is used to make pseudo labels of target domain, and further extracted the text embeddings and image embeddings of source domain and target domain. And then they used image embedding to guide the backbone network to learn the self-enhanced features and cross-domain features. Different from PDA (Bai et al., 2024), we achieved better classification results which benefits from injecting the semantic information into backbone network. Compared with PDA (Bai et al., 2024), which also utilizes VLMs, our proposed method performs more than 7.42 % on the classification accuracy.
Table 3. Ablation study for classification accuracy on RMO
@@
61.614
67.253
69.385
72.287
73.315
4.5. Ablation Study
@@
We combine ViT-B/16 (Dosovitskiy et al., 2021) with adversarial training as the baseline and obtain a 61.612% on the classification accuracy of RMO. Adding offsets on input images directly makes a 5.639% increase compared to baseline in classification accuracy. Comparison of the second and third rows in Table 3, replacing the direct offset on images with adding random offset to the token sequence further improves the classification accuracy by 2.132%. It proves that injecting perturbations into the hidden space is a better choice. On this basis, the Generalization Enhancement Module is introduced to make the classification accuracy finally rise to 72.287%, which shows that it is meaningful to use CLIP for Self-Knowledge Distillation on the source domain. Finally, we employ curriculum learning to regularly select the source domain samples for training, which further improves the classification accuracy to 73.315 %. The result demonstrates the effectiveness of Adaptive Adjustment Mechanism. In summary, the ablation studies conducted on the RMO effectively validate the contributions of the proposed modules.
4.6. Parameter Sensitivity Discussion
@@
In this part, we analyze the sensitivity of three key parameters in our framework: $\alpha$ and $\beta$ used in Eq. 3, and $\gamma$ in Eq. 5. The classification accuracy trends under varying parameters are visualized in Fig. 4. As shown in Fig. 4(a), the stable variation validates the robustness of the proposed Generalization Enhancement module. Fig. 4(b) reveals a critical dependency on $\gamma$, which underscores that the necessity of sufficient regularization that overly small $\gamma$ fails to mitigate overfitting, while larger values retain competitive accuracy. And the results in Fig. 4(c) imply that insufficient weight weaken its correction effect, while overemphasis may distort the primary learning objective. Therefore, the best choices for these parameters are $\alpha=0.3$, $\beta=1$, and $\gamma=0.2$.
Figure 4. Parameter sensitivity experiments.
5. Conclusion
@@
In this work, we propose a label-rich and balanced AI-generated Benchmark for Maritime Object Classification (AIMO) with multiple weather and illumination conditions. And we construct a Real Maritime Object Benchmark (RMO) collected from a series of real maritime object datasets and the Google. We further introduce an UDA framework with adversarial training and curriculum learning based on Vision Transformer to address maritime object classification in the real-world scenarios. It leverages the generalization of VLM to refine source domain features and narrow domain gap. Experiments validate the effectiveness of the proposed benchmark and unsupervised domain adaptation framework.

### v3

@@ 4.4 Effectiveness of Domain Adaptation
Comparison with the basic networks. ViT-small DBLP:conf/iclr/DosovitskiyB0WZ21 and ViT-base DBLP:conf/iclr/DosovitskiyB0WZ21 have applied Transformer to image classification for the first time, demonstrating the potential of global attention in visual tasks. On this basis, Swin Transformer liu2021swin takes into account global modeling and local concentration bias through windowing and hierarchical design. Compared with these basic networks, our proposed method gained of over 10 % on the classification accuracy, which attributes to our effective designs based on a series of modules in order to mitigate domain divergence.
@@
Comparison with methods using pseudo labels. CDTrans xu2021cdtrans learned domain-invariant features through cross-domain attention and used target domain pseudo labels for self-supervised learning. Given the substantial differences in distribution between AIMO and RMO datasets, directly aligning image pairs from the source and target domains is not an effective approach. And the large domain gap makes the pseudo labels more unreliable, which directly affects the model performance. Therefore, CDTrans xu2021cdtrans exhibits suboptimal performance. Also, different from PMTrans zhu2023patch, which constructed intermediate domains to alleviate domain gap and also relied on the pseudo labels during training, our method utilizes VLMs to alleviate the domain-shift from source domain to target domain, achieving nearly 8.2 % gains on the accuracy.
@@
Comparison with method using adversarial adaptation. TVT yang2023tvt designed a transferability adaption module based on adversarial adaptation. The learned transferabilities were injected into attention blocks, promoting ViT to focus on both domain-invariant and domain-specific features. Compared with TVT yang2023tvt, which also utilizes adversarial training for UDA, the accuracy of our proposed method is 11 % higher, probably due to the use of VLMs (i.e., CLIP) for improving robustness of the source domain features.
Comparison with methods using VLMs. PDA bai2024prompt utilized the zero-shot capability of CLIP to make pseudo labels of target domain, and further extracted the text embeddings and image embeddings of source domain and target domain. UniMoS li2024split was a Unified Modality Separation framework for unsupervised domain adaptation, which distinctly disentangles CLIP’s features into language-associated and vision-associated components. DAMP du2024domain mutually aligned visual and textual embeddings based on a cross-attention module, a semantic-consistency loss and an instance-discrimination contrastive loss. Different from the methods above, we achieved better classification results which benefits from injecting the semantic information into backbone network through self-knowledge distillation loss. Also, we utilize curriculum learning strategy to reduce data overfitting, which further improves performance.
Comparison with methods on Mini-DomainNet. To demonstrate the generality of our proposed method in UDA, we conduct comparative experiments with other methods on the Mini-DomainNet dataset, which is a subset of the most challenging dataset DomainNet DBLP:conf/iccv/PengBXHSW19 in Table 3. The results proved the generalization of our proposed method.
Table 4: Ablation study for classification accuracy on RMO
@@
61.629
67.416
69.443
71.337
73.675
Table 5: Classification accuracy on RMO with different prompt template
Prompt Template Acc (%)
A photo in {domain} 69.570
A photo of a {class} 71.979
A photo of a {class} in {domain} 73.675
4.5 Ablation Study
@@
We combine ViT-B/16 DBLP:conf/iclr/DosovitskiyB0WZ21 with adversarial training as the baseline and obtain a 61.629% on the classification accuracy of RMO. Adding offsets on input images directly makes a 5.787% increase compared to baseline in classification accuracy. Comparison of the second and third rows in Table 4, replacing the direct offset on images with adding random offset to the token sequence further improves the classification accuracy by 2.027%. It proves that injecting perturbations into the hidden space is a better choice. On this basis, the Generalization Enhancement Module is introduced to make the classification accuracy finally rise to 71.337%, which shows that it is meaningful to use CLIP for Self-Knowledge Distillation on the source domain. Finally, we employ curriculum learning to regularly select the source domain samples for training, which further improves the classification accuracy to 73.675 %. The result demonstrates the effectiveness of Adaptive Adjustment Mechanism. In summary, the ablation studies conducted on the RMO effectively validate the contributions of the proposed modules.
Moreover, we further compare the selection of prompt for text embedddings. As shown in Table 5, when we use "A photo of a {class} in {domain}" as prompt template, the classification accuracy achieved the highest value. This is due to the comprehensive consideration of the influence of category and weather on the classification accuracy. Compared with using "{class}" or "{domain}" alone, this compositive prompt template allows the generalization enhancement module to contain more information about the image.
Figure 3: Validation experiments with our proposed method and SCAN. (a): Comparison of accuracy on RMO with multiple weather & illumination conditions. (b): Comparison of accuracy on RMO with different categories.
Figure 4: Parameter sensitivity experiments.
4.6 Parameter Sensitivity Discussion
@@
In this part, we analyze the sensitivity of three key parameters in our framework: $\alpha$ and $\beta$ used in Eq. 4, and $\gamma$ in Eq. 6. The classification accuracy trends under varying parameters are visualized in Fig. 4. As shown in Fig. 4(a), the model maintains stable performance across $\alpha\in[0.1,0.4]$, with accuracy fluctuating narrowly between 68.56% and 73.68%. This minimal variation validates the robustness of the proposed Generalization Enhancement module. We obverse that the choice of $\alpha$ has little impact on the performance of the model, which proves the stability of our proposed Generalization Enhancement Module. Fig. 4(b) reveals a critical dependency on $\gamma$, because $\gamma$ influences the strength of regularization. When $\gamma=0.2$, the model achieves peak accuracy, but performance plummets a 9.98% degradation at $\gamma=0.1$. This sharp decline underscores the necessity of sufficient regularization that overly small $\gamma$ fails to mitigate overfitting, while larger values retain competitive accuracy. While, when the value of $\gamma$ is too small, the classification accuracy drops significantly, because the regularization effect is not achieved. Similarly, the weight of the refinement loss, $\beta$, also has an impact on the classification accuracy of the model. The classifier refinement loss weight $\beta$ exhibits a clear optimum at $\beta=1$. Deviations from this value, degrade performance by 3.9–7.6%, as shown in Fig. 4(c). This implies that insufficient weight weaken its correction effect, while overemphasis may distort the primary learning objective. Therefore, the best choices for these parameters are $\alpha=0.3$, $\beta=1$, and $\gamma=0.2$.
5 Conclusion
@@
In this work, we propose a label-rich and balanced AI-generated Benchmark for Maritime Object Classification (AIMO) with multiple weather and illumination conditions. And we construct a Real Maritime Object Benchmark (RMO) collected from a series of real maritime object datasets and the Google. We further introduce an UDA framework with adversarial training and curriculum learning based on Vision Transformer to address maritime object classification in the real-world scenarios. It leverages the generalization of VLM to refine source domain features and narrow domain gap. And an adaptive adjustment mechanism is applied to prevent model collapse and data overfitting. Experiments validate the effectiveness of the proposed benchmark and unsupervised domain adaptation framework.
[…]